Sol Summary – May 22nd

Crew 281 Sol Summary Report 22 May 2023
Sol: 8
Summary Title: Hair wash day
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
Today’s EVA was at Pooh’s corner. The crew of 3 drove Spirit and Perseverance rovers and reached the site at 0823 hrs. They captured scanline measurements for the boulders Ana Pires (Crew Scientist) approved at Pooh’s Corner. KC Shasteen helped with the Equotip measurements while Ana gathered the Schmidt hammer data. The crew finished a 5 station long scanline and during their measurements Ritupriya Patil (XO, EVA Lead) deployed the drone for aerial imagery. Detailed vertical images and videos of the scanline were captured to help confirm the weathering grade, discontinuities of the rocks. These images and videos will be used for georeferencing the area around the scanline and for further assessment of the geological region. The objective was achieved right before winds gusted to 27 kmph and the drone was secured.
For lunch Ritu (XO) made tofu Thai curry with white rice. Post lunchtime Ritu and Ana worked on tomorrow’s EVA planning, looking at maps, distance and noting time for each stop over during the 2 person EVA. Then Ritu and KC worked in the RAM to build a drone packaging so it can be securely stored during the travel time in a ready to be deployed configuration. This makes the drone launch much faster and efficient.
Rachel (HSO) checked into OMISS Old Man International Side band Society through the ham radio setup. She is working through troubleshooting the power reception she is having. While Ana worked on closing the set of scanlines and working further on research and synergies with Ritu. KC and Rachel had a hair wash day to clear their heads (get that?!).
In the green hab, Megan (Commander) watered and checked on the plants. KC (Crew Engineer) did routine checks on hab operations – batteries, tunnels, flushed toilet and water levels. Ana cleaned up her robotic scoop in the RAM after a successful mission today.
In the science dome Megan checked on the cacao seeds and made a video recording the changes to the seeds.
For dinner Rachel made stamppotten a mashed potato dish from the Netherland. It rained in the late afternoon and there were storms, but the antenna stayed upright.
Look Ahead Plan: Tomorrow we will do a 2-person, long range EVA to the Overlook region. It will be an all-female crew exploring the Martian surface and capturing aerial footage. On our way back we will stop at Pooh’s Corner for images of scanlines that were measured on earlier EVAs.
Anomalies in work: None
Weather: Windy
Crew Physical Status: Nominal
EVA: Completed 1 today

Reports to be filed:
Greenhab Report
EVA Report
EVA Request
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms if the antenna needs to come down or be retrieved.

Copyright © The Mars Society. All rights reserved. | Main Site